according to official censes figures,8% of couples living together are not arried.A researcher took a random sample of 400 couples and found that 9.5% of them are not married.Test at the 15% significanve level if the current percentage of unmarried couples is different from 8%


1
Expert's answer
2021-08-04T16:49:57-0400

"H_0:\pi=0.08."

"H_a:\pi\ne0.08."

Tests statistic: "z=\frac{0.095-0.08}{\sqrt{\frac{0.08(1-0.08)}{400}}}=1.11."

P-value: "p=2P(Z>1.11)=0.2670."

Since the p-value is greater than 0.15, fail to reject the null hypothesis.

There is no sufficient evidence that the current percentage of unmarried couples is different from 8%.